# Bias and Fairness in Large Language Models: A Survey

### v2

@@ 2.2.3 Bias in the Development and Deployment Life Cycle
Another way of understanding social bias in LLMs is to examine at which points within the model development and deployment process the bias emerges, which may exacerbate preexisting historical biases. This has been thoroughly explored by Mehrabi et al. (2021) and Suresh & Guttag (2021), and we summarize these pathways here:
@@
  • Deployment: An LLM may be deployed in a different setting than that for which it was intended, such as with or without a human intermediary for automated decision-making. The interface through which a user interacts with the model, such as the presentation of model outputs, may change human perception of the LLM’s behavior.
@@ 2.3 Fairness Desiderata for LLMs
Though group, individual, and subgroup fairness define useful general frameworks, they in themselves do not specify the exact fairness constraints. This distinction is critical, as defining the "right" fairness specification is highly subjective, value-dependent, and non-static, evolving through time (Barocas et al., 2019; Ferrara, 2023; Friedler et al., 2021). Each stakeholder brings perspectives that may specify different fairness constraints for the same application and setting. In sum, the is no universal fairness specification.
Instead of suggesting a single fairness constraint, we provide a number of possible fairness desiderata for LLMs. While similar concepts have been operationalized for machine learning classification tasks (Mehrabi et al., 2021; Verma & Rubin, 2018), less has been done in the NLP space, which may contain more ambiguity than classification for tasks like language generation. Note that for NLP classification tasks, or tasks with a superimposed classifier, traditional fairness definitions like equalized odds or statistical parity may be used without modification. For cases when simple classification may not be useful, we present general desiderata of fairness for NLP tasks that generalize notions in the LLM bias evaluation and mitigation literature. We use the following notation: for some input $X_{i}$ containing a mention of a social group $G_{i}$, let $X_{j}$ be an analogous input with the social group substituted for $G_{j}$. Let $w\in W$ be a neutral word, and let $a\in A$ be a protected attribute word, with $a_{i}$ and $a_{j}$ as corresponding terms associated with $G_{i}$ and $G_{j}$, respectively. Let $X_{\setminus A}$ represent an input with all social group identifiers removed. See Table 2 for this and other notation.
@@ 2.4.1 Taxonomy of Metrics for Bias Evaluation
We summarize several evaluation techniques that leverage a range of fairness desiderata and operate at different fundamental levels. As the subset of appropriate evaluation metrics $\psi(\mathcal{D})\subseteq\Psi$ is largely determined by (i) access to the model (i.e., access to trainable model parameters, versus access to model output only) and (ii) the data structure of an evaluation set $\mathcal{D}$, we taxonomize metrics by the underlying data structure assumed by the metric. The complete taxonomy is described in Section 3.
@@
      Sentence Embedding (§3.3.2): Adapt word embedding metrics to contextualized embeddings
@@
      Masked Token (§3.4.1): Compare token probabilities from fill-in-the-blank templates
@@ 2.4.2 Taxonomy of Datasets for Bias Evaluation
Bias evaluation datasets can test an LLM for specific harms, such as stereotyping or derogatory language, that target particular social groups, such as gender or race groups. Similar to our taxonomy of metrics, we organize datasets by their data structure. The complete taxonomy is described in Section 4.
@@
    Counterfactual Inputs: Compare pairs or tuples of sentences with perturbed social groups
@@
      Masked Tokens (§4.1.1): Ask an LLM to predict the most likely word via fill-in-the-blank
@@
      Unmasked Sentences (§4.1.2): Ask an LLM to predict the most likely sentence
@@
      Sentence Completions (§4.2.1): Ask an LLM to provide a continuation
@@
      Question-Answering (§4.2.2): Ask an LLM to select an answer to a question
@@ 2.4.3 Taxonomy of Techniques for Bias Mitigation
      Data Augmentation (§5.1.1): Add new training data to extend the distribution
@@
      Data Filtering and Reweighting (§5.1.2): Remove or reweight training instances
@@
      Data Generation (§5.1.3): Produce a new dataset meeting a set of standards or characteristics
@@
      Instruction Tuning (§5.1.4): Prepend additional tokens to an input to condition the output
@@
      Projection-based Mitigation (§5.1.5): Transform representations in the embedding space
@@
      Loss Function Modification (§5.2.2): Introduce a new equalizing objective function
@@
      Selective Parameter Updating (§5.2.3): Fine-tune a subset of model parameters
@@
      Filtering Model Parameters (§5.2.4): Remove a subset of model parameters
@@
      Decoding Strategy Modification (§5.3.1): Modify next-word or sequence probabilities
@@
      Weight Redistribution (§5.3.2): Modify the entropy of the LLM’s attention weights
@@
      Modular Debiasing Networks (§5.3.3): Create stand-alone debiasing components
@@ 3.1 Facets of Evaluation of Biases: Metrics and Datasets
Figure 1: Fairness Evaluation via Substitution. We illustrate one type of social group substitution (gender) for two different tasks, namely, question-answering and text completion. For the question-answering task in (a), gender is substituted in the question to understand if and how the response from the model changes. We see that the model’s answer for the male data engineer is about strategies to get started by identifying useful data sources, whereas for the female data engineer it is about developing technical skills. There is an implicit assumption that male data engineers already have the technical skills they need compared to female data engineers. For the completion task in (b), we again substitute the gender, and see that this specific model responds the same, that is, it completes the sentence by generating the same text for either case. While in both cases we fix the profession, this is another factor that can be varied. Other types of bias can also be evaluated using different metrics, such as toxicity or sentiment. We also note that in (a) the gender is more explicit compared to (b) where it is more implicit. This is another dimension that can be evaluated.
@@
However, other pseudo-log-likelihood-based metrics can also be computed from the CrowS-Pairs dataset, such as All Unmasked Likelihood (Kaneko & Bollegala, 2022). Therefore, whenever possible, we decompose the dataset from the metric that was originally used over it. In our taxonomy of datasets in Section 4, we discuss potential alternative metrics that can be used with various classes of datasets.
@@ 3.2 Taxonomy of Metrics based on What They Use
$(a_{1},\cdots,a_{m})$ protected attributes words with analogous meanings for groups $G_{1},\cdots,G_{m}$
@@
$\bm{\mathrm{g}}$ gender direction in embedding space
$B$ gender subspace in embedding space
@@
$E(\cdot)$ word or sentence encoder
@@
Table 3: Taxonomy of Evaluation Metrics for Bias Evaluation in LLMs. We summarize metrics that measure bias using embeddings, model-assigned probabilities, or generated text. The data structure describes the input to the model required to compute the metrics, and $\mathcal{D}$ indicates if the metric was introduced with an accompanying dataset. $W$ is the set of neutral words; $A_{i}$ is the set of sensitive attribute words associated with group $G_{i}$; $S\in\mathbb{S}$ is a (masked) input sentence or template, which may be neutral ($S_{W}$) or contain sensitive attributes ($S_{A}$); $M$ and $U$ are the sets of masked and unmasked tokens in $S$, respectively; $\hat{Y}_{i}\in\hat{\mathbb{Y}}$ is a predicted output associated with group $G_{i}$; $c(\cdot)$ is a classifier; $PP(\cdot)$ is perplexity; $\psi(\cdot)$ is an invariance metric; $C(\cdot)$ is a co-occurrence count; and $\mathcal{W}_{1}(\cdot)$ is Wasserstein-1 distance. For additional notation, see Table 2.
Metric Data Structure Equation $\mathcal{D}$
@@
Word Embedding* (§ 3.3.1)
  $\textrm{}^{\dagger}$ Static word
$f(A,W)=({\textrm{mean}}_{a_{1}\in A_{1}}s(a_{1},W_{1},W_{2})$
$\times$
  $-{\textrm{mean}}_{a_{2}\in A_{2}}s(a_{2},W_{1},W_{2}))/{\textrm{std}}_{a\in A}s(a,W_{1},W_{2})$
$\times$
@@
  SEAT Contextual sentence
$f(S_{A},S_{W})=\textrm{WEAT}(S_{A},S_{W})$
$\times$
  CEAT Contextual sentence
$f(S_{A},S_{W})=\tfrac{\Sigma_{i=1}^{N}v_{i}\textrm{WEAT}(S_{A_{i}},S_{W_{i}})}{\Sigma_{i=1}^{N}v_{i}}$
$\times$
  Sentence Bias Score Contextual sentence
$f(S)=\sum_{s\in S}|\cos(\bm{\mathrm{s}},\bm{\mathrm{g}})\cdot\alpha_{s}|$
$\checkmark$
@@
  DisCo Masked
$f(S)=\mathbb{I}(\hat{y}_{i,\texttt{[MASK]}}=\hat{y}_{j,\texttt{[MASK]}})$
$\times$
  Log-Probability Bias Score Masked
$f(S)=\log\tfrac{p_{a_{i}}}{p_{prior_{i}}}-\log\tfrac{p_{a_{j}}}{p_{prior_{j}}}$
$\times$
  Categorical Bias Score Masked
$f(S)=\tfrac{1}{|W|}\Sigma_{w\in W}\textrm{Var}_{a\in A}\log\tfrac{p_{a}}{p_{prior}}$
$\times$
Pseudo-Log-Likelihood (§ 3.4.2)
$f(S)=\mathbb{I}(g(S_{1})>g(S_{2}))$
  CrowS-Pairs Score Stereo, anti-stereo
$g(S)=\Sigma_{u\in U}\log P(u|U_{\setminus u},M;\theta)$
$\checkmark$
  Context Association Test Stereo, anti-stereo
$g(S)=\tfrac{1}{|M|}\Sigma_{m\in M}\log P(m|U;\theta)$
$\checkmark$
  All Unmasked Likelihood Stereo, anti-stereo
$g(S)=\tfrac{1}{|S|}\Sigma_{s\in S}\log P(s|S;\theta)$
$\times$
  Language Model Bias Stereo, anti-stereo
$f(S)=t\textrm{-value}(PP(S_{1}),PP(S_{2}))$
$\checkmark$
@@
  Social Group Substitution Counterfactual pair
$f(\hat{Y})=\psi(\hat{Y}_{i},\hat{Y}_{j})$
$\times$
  Co-Occurrence Bias Score Any prompt
$f(w)=\log\tfrac{P(w|A_{i})}{P(w|A_{j})}$
$\times$
  Demographic Representation Any prompt
$f(G)=\Sigma_{a\in A}\Sigma_{\hat{Y}\in\hat{\mathbb{Y}}}C(a,\hat{Y})$
$\times$
  Stereotypical Associations Any prompt
$f(w)=\Sigma_{a\in A}\Sigma_{\hat{Y}\in\hat{\mathbb{Y}}}C(a,\hat{Y})\mathbb{I}(C(w,\hat{Y})>0)$
$\times$
@@
  Perspective API Toxicity prompt
$f(\hat{Y})=c(\hat{Y})$
$\times$
   Expected Maximum Toxicity Toxicity prompt
$f(\hat{\mathbb{Y}})=\text{max}_{\hat{Y}\in\hat{\mathbb{Y}}}c(\hat{Y})$
$\times$
   Toxicity Probability Toxicity prompt
$f(\hat{\mathbb{Y}})=P(\sum_{\hat{Y}\in\hat{\mathbb{Y}}}\mathbb{I}(c(\hat{Y})\geq 0.5)\geq 1)$
$\times$
   Toxicity Fraction Toxicity prompt
$f(\hat{\mathbb{Y}})=\mathbb{E}_{\hat{Y}\in\hat{\mathbb{Y}}}[\mathbb{I}(c(\hat{Y})\geq 0.5)]$
$\times$
  Score Parity Counterfactual pair
$f(\hat{\mathbb{Y}})=|\mathbb{E}_{\hat{Y}\in\hat{\mathbb{Y}}}[c(\hat{Y}_{i},i)|A=i]-\mathbb{E}_{\hat{Y}\in\hat{\mathbb{Y}}}[c(\hat{Y}_{j},j)|A=j]|$
$\times$
  Counterfactual Sentiment Bias Counterfactual pair
$f(\hat{\mathbb{Y}})=\mathcal{W}_{1}(P(c(\hat{\mathbb{Y}}_{i})|A=i),P(c(\hat{\mathbb{Y}}_{j}|A=j))$
$\times$
  Regard Score Counterfactual tuple
$f(\hat{Y})=c(\hat{Y})$
$\times$
  Full Gen Bias Counterfactual tuple
$f(\hat{\mathbb{Y}})=\Sigma_{i=1}^{C}\textrm{Var}_{w\in W}(\frac{1}{|\hat{\mathbb{Y}}_{w}|}\Sigma_{\hat{Y}_{w}\in\hat{\mathbb{Y}}_{w}}c(\hat{Y}_{w})[i])$
$\checkmark$
@@
  HONEST Counterfactual tuple
$f(\hat{\mathbb{Y}})=\tfrac{\Sigma_{\hat{Y}_{k}\in\hat{\mathbb{Y}}_{k}}\Sigma_{\hat{y}\in\hat{Y}_{k}}\mathbb{I}_{\textrm{HurtLex}}(\hat{y})}{|\hat{\mathbb{Y}}|\cdot k}$
$\times$
  Psycholinguistic Norms Any prompt
$f(\hat{\mathbb{Y}})=\tfrac{\Sigma_{\hat{Y}\in\hat{\mathbb{Y}}}\Sigma_{\hat{y}\in\hat{Y}}\textrm{sign}(\textrm{affect-score}(\hat{y}))\textrm{affect-score}(\hat{y})^{2}}{\Sigma_{\hat{Y}\in\hat{\mathbb{Y}}}\sum_{\hat{y}\in\hat{Y}}|\textrm{affect-score}(\hat{y})|}$
$\checkmark$
  Gender Polarity Any prompt
$f(\hat{\mathbb{Y}})=\tfrac{\Sigma_{\hat{Y}\in\hat{\mathbb{Y}}}\Sigma_{\hat{y}\in\hat{Y}}\textrm{sign}(\textrm{bias-score}(\hat{y}))\textrm{bias-score}(\hat{y})^{2}}{\Sigma_{\hat{Y}\in\hat{\mathbb{Y}}}\sum_{\hat{y}\in\hat{Y}}|\textrm{bias-score}(\hat{y})|}$
$\checkmark$
*Static word embeddings are not used with LLMs, but we include the word embedding metric WEAT for completeness given its relevance to sentence embedding metrics. $\textrm{}^{\dagger}$ee § 3.3.1 for definition of $s(\cdot)$.

### v3

@@ 2.2.3 Bias in the Development and Deployment Life Cycle
Another way of understanding social bias in LLMs is to examine at which points within the model development and deployment process the bias emerges, which may exacerbate preexisting historical biases. This has been thoroughly explored by Mehrabi et al. (2021), Shah, Schwartz, and Hovy (2020), and Suresh and Guttag (2021), and we summarize these pathways here:
@@
  • Deployment: An LLM may be deployed in a different setting than that for which it was intended, such as with or without a human intermediary for automated decision-making. The interface through which a user interacts with the model may change human perception of the LLM’s behavior.
@@ 2.3 Fairness Desiderata for LLMs
Though group, individual, and subgroup fairness define useful general frameworks, they in themselves do not specify the exact fairness constraints. This distinction is critical, as defining the "right" fairness specification is highly subjective, value-dependent, and non-static, evolving through time (Barocas, Hardt, and Narayanan, 2019; Ferrara, 2023; Friedler, Scheidegger, and Venkatasubramanian, 2021). Each stakeholder brings perspectives that may specify different fairness constraints for the same application and setting. The list — and the accompanying interests — of stakeholders is broad. In the machine learning data domain more broadly, Jernite et al. (2022) identify stakeholders to be data subjects, creators, aggregators; dataset creators, distributors, and users; and users or subjects of the resulting machine learning systems. Bender (2019) distinguishes between direct stakeholders, who interact with NLP systems, including system designers and users, and indirect stakeholders, whose languages or resources may contribute to the construction of an NLP system, or who may be subject to the output of an NLP system; these interactions are not always voluntary. In sum, the is no universal fairness specification.
Instead of suggesting a single fairness constraint, we provide a number of possible fairness desiderata for LLMs. While similar concepts have been operationalized for machine learning classification tasks (Mehrabi et al., 2021; Verma and Rubin, 2018), less has been done in the NLP space, which may contain more ambiguity than classification for tasks like language generation. Note that for NLP classification tasks, or tasks with a superimposed classifier, traditional fairness definitions like equalized odds or statistical parity may be used without modification. For cases when simple classification may not be useful, we present general desiderata of fairness for NLP tasks that generalize notions in the LLM bias evaluation and mitigation literature, building on the outcome and error disparity definitions proposed by Shah, Schwartz, and Hovy (2020). We use the following notation: for some input $X_{i}$ containing a mention of a social group $G_{i}$, let $X_{j}$ be an analogous input with the social group substituted for $G_{j}$. Let $w\in W$ be a neutral word, and let $a\in A$ be a protected attribute word, with $a_{i}$ and $a_{j}$ as corresponding terms associated with $G_{i}$ and $G_{j}$, respectively. Let $X_{\setminus A}$ represent an input with all social group identifiers removed. See Table 2 for this and other notation.
@@ 2.4.1 Taxonomy of Metrics for Bias Evaluation
We summarize several evaluation techniques that leverage a range of fairness desiderata and operate at different fundamental levels. As the subset of appropriate evaluation metrics $\psi(\mathcal{D})\subseteq\Psi$ is largely determined by (1) access to the model (i.e., access to trainable model parameters, versus access to model output only) and (2) the data structure of an evaluation set $\mathcal{D}$, we taxonomize metrics by the underlying data structure assumed by the metric. The complete taxonomy is described in Section 3.
@@
      Sentence Embedding (§3.3.2): Adapt to contextualized embeddings
@@
      Masked Token (§3.4.1): Compare fill-in-the-blank probabilities
@@ 2.4.2 Taxonomy of Datasets for Bias Evaluation
Bias evaluation datasets can assess specific harms, such as stereotyping or derogatory language, that target particular social groups, such as gender or race groups. Similar to our taxonomy of metrics, we organize datasets by their data structure. The complete taxonomy is described in Section 4.
@@
    Counterfactual Inputs: Compare sets of sentences with perturbed social groups
@@
      Masked Tokens (§4.1.1): LLM predicts the most likely fill-in-the-blank
@@
      Unmasked Sentences (§4.1.2): LLM predicts the most likely sentence
@@
      Sentence Completions (§4.2.1): LLM provides a continuation
@@
      Question-Answering (§4.2.2): LLM selects an answer to a question
@@ 2.4.3 Taxonomy of Techniques for Bias Mitigation
      Data Augmentation (§5.1.1): Extend distribution with new data
@@
      Data Filtering and Reweighting (§5.1.2): Remove or reweight instances
@@
      Data Generation (§5.1.3): Produce new data meeting certain standards
@@
      Instruction Tuning (§5.1.4): Prepend additional tokens to an input
@@
      Projection-based Mitigation (§5.1.5): Transform hidden representations
@@
      Loss Function Modification (§5.2.2): Introduce a new objective
@@
      Selective Parameter Updating (§5.2.3): Fine-tune a subset of parameters
@@
      Filtering Model Parameters (§5.2.4): Remove a subset of parameters
@@
      Decoding Strategy Modification (§5.3.1): Modify probabilities
@@
      Weight Redistribution (§5.3.2): Modify the entropy of attention weights
@@
      Modular Debiasing Networks (§5.3.3): Add stand-alone components
@@ 3.1 Facets of Evaluation of Biases: Metrics and Datasets
Figure 1: Evaluation via Substitution for Two Tasks. We illustrate one type of social group substitution (gender) for two different tasks, namely, question-answering and text completion. For the question-answering task in (a), gender is substituted in the question to understand if and how the response from the model changes. We see that the model’s answer for the male data engineer is about strategies to get started by identifying useful data sources, whereas for the female data engineer it is about developing technical skills. There is an implicit assumption that male data engineers already have the technical skills they need compared to female data engineers. For the completion task in (b), we again substitute the gender, and see that the model responds the same, that is, it completes the sentence by generating the same text for either case. We note that in (a) the gender is more explicit compared to (b) where it is more implicit.
@@
However, other pseudo-log-likelihood-based metrics can also be computed from the CrowS-Pairs dataset, such as All Unmasked Likelihood (Kaneko and Bollegala, 2022). Therefore, whenever possible, we decompose the dataset from the metric that was originally used over it. In our taxonomy of datasets in Section 4, we discuss potential alternative metrics that can be used with various classes of datasets.
@@ 3.2 Taxonomy of Metrics based on What They Use
$(a_{1},\cdots,a_{m})$ protected attributes with analogous meanings for $G_{1},\cdots,G_{m}$
@@
$\bm{\mathrm{v}}_{\textrm{gender}}$ gender direction in embedding space
$V_{\textrm{gender}}$ gender subspace in embedding space
@@
$E(\cdot)$ word or sentence embedding
@@
Table 3: Taxonomy of Evaluation Metrics for Bias Evaluation in LLMs. We summarize metrics that measure bias using embeddings, model-assigned probabilities, or generated text. The data structure describes the input to the model required to compute the metrics, and $\mathcal{D}$ indicates if the metric was introduced with an accompanying dataset. $W$ is the set of neutral words; $A_{i}$ is the set of sensitive attribute words associated with group $G_{i}$; $S\in\mathbb{S}$ is a (masked) input sentence or template, which may be neutral ($S_{W}$) or contain sensitive attributes ($S_{A}$); $M$ and $U$ are the sets of masked and unmasked tokens in $S$, respectively; $\hat{Y}_{i}\in\hat{\mathbb{Y}}$ is a predicted output associated with group $G_{i}$; $c(\cdot)$ is a classifier; $PP(\cdot)$ is perplexity; $\psi(\cdot)$ is an invariance metric; $C(\cdot)$ is a co-occurrence count; $\mathcal{W}_{1}(\cdot)$ is Wasserstein-1 distance; and $\mathbb{E}$ is the expected value.
Metric Data Structure* Equation $\mathcal{D}$
@@
Word $\textrm{}^{\dagger}$ (§ 3.3.1)
   $\textrm{}^{\ddagger}$ Static word $f(A,W)=({\textrm{mean}}_{a_{1}\in A_{1}}s(a_{1},W_{1},W_{2})$$\times$
  $-{\textrm{mean}}_{a_{2}\in A_{2}}s(a_{2},W_{1},W_{2}))/{\textrm{std}}_{a\in A}s(a,W_{1},W_{2})$$\times$
@@
   SEAT Contextual sentence $f(S_{A},S_{W})=\textrm{WEAT}(S_{A},S_{W})$$\times$
   CEAT Contextual sentence $f(S_{A},S_{W})=\tfrac{\Sigma_{i=1}^{N}v_{i}\textrm{WEAT}(S_{A_{i}},S_{W_{i}})}{\Sigma_{i=1}^{N}v_{i}}$$\times$
   Sentence Bias Score Contextual sentence $f(S)=\sum_{s\in S}|\cos(\bm{\mathrm{s}},\bm{\mathrm{v}}_{\textrm{gender}})\cdot\alpha_{s}|$checkmark
@@
   DisCo Masked $f(S)=\mathbb{I}(\hat{y}_{i,\texttt{[MASK]}}=\hat{y}_{j,\texttt{[MASK]}})$$\times$
   Log-Probability Bias Score Masked $f(S)=\log\tfrac{p_{a_{i}}}{p_{prior_{i}}}-\log\tfrac{p_{a_{j}}}{p_{prior_{j}}}$$\times$
   Categorical Bias Score Masked $f(S)=\tfrac{1}{|W|}\Sigma_{w\in W}\textrm{Var}_{a\in A}\log\tfrac{p_{a}}{p_{prior}}$$\times$
Pseudo-Log-Likelihood (§ 3.4.2) $f(S)=\mathbb{I}(g(S_{1})>g(S_{2}))$
   CrowS-Pairs Score Stereo, anti-stereo $g(S)=\Sigma_{u\in U}\log P(u|U_{\setminus u},M;\theta)$checkmark
   Context Association Test Stereo, anti-stereo $g(S)=\tfrac{1}{|M|}\Sigma_{m\in M}\log P(m|U;\theta)$checkmark
   All Unmasked Likelihood Stereo, anti-stereo $g(S)=\tfrac{1}{|S|}\Sigma_{s\in S}\log P(s|S;\theta)$$\times$
   Language Model Bias Stereo, anti-stereo $f(S)=t\textrm{-value}(PP(S_{1}),PP(S_{2}))$checkmark
@@
   Social Group Substitution Counterfactual pair $f(\hat{Y})=\psi(\hat{Y}_{i},\hat{Y}_{j})$$\times$
   Co-Occurrence Bias Score Any prompt $f(w)=\log\tfrac{P(w|A_{i})}{P(w|A_{j})}$$\times$
   Demographic Representation Any prompt $f(G)=\Sigma_{a\in A}\Sigma_{\hat{Y}\in\hat{\mathbb{Y}}}C(a,\hat{Y})$$\times$
   Stereotypical Associations Any prompt $f(w)=\Sigma_{a\in A}\Sigma_{\hat{Y}\in\hat{\mathbb{Y}}}C(a,\hat{Y})\mathbb{I}(C(w,\hat{Y})>0)$$\times$
@@
   Perspective API Toxicity prompt $f(\hat{Y})=c(\hat{Y})$$\times$
    Expected Maximum Toxicity Toxicity prompt $f(\hat{\mathbb{Y}})=\text{max}_{\hat{Y}\in\hat{\mathbb{Y}}}c(\hat{Y})$$\times$
    Toxicity Probability Toxicity prompt $f(\hat{\mathbb{Y}})=P(\sum_{\hat{Y}\in\hat{\mathbb{Y}}}\mathbb{I}(c(\hat{Y})\geq 0.5)\geq 1)$$\times$
    Toxicity Fraction Toxicity prompt $f(\hat{\mathbb{Y}})=\mathbb{E}_{\hat{Y}\in\hat{\mathbb{Y}}}[\mathbb{I}(c(\hat{Y})\geq 0.5)]$$\times$
   Score Parity Counterfactual pair $f(\hat{\mathbb{Y}})=|\mathbb{E}_{\hat{Y}\in\hat{\mathbb{Y}}}[c(\hat{Y}_{i},i)|A=i]-\mathbb{E}_{\hat{Y}\in\hat{\mathbb{Y}}}[c(\hat{Y}_{j},j)|A=j]|$$\times$
   Counterfactual Sentiment Bias Counterfactual pair $f(\hat{\mathbb{Y}})=\mathcal{W}_{1}(P(c(\hat{\mathbb{Y}}_{i})|A=i),P(c(\hat{\mathbb{Y}}_{j}|A=j))$$\times$
   Regard Score Counterfactual tuple $f(\hat{Y})=c(\hat{Y})$$\times$
   Full Gen Bias Counterfactual tuple $f(\hat{\mathbb{Y}})=\Sigma_{i=1}^{C}\textrm{Var}_{w\in W}(\frac{1}{|\hat{\mathbb{Y}}_{w}|}\Sigma_{\hat{Y}_{w}\in\hat{\mathbb{Y}}_{w}}c(\hat{Y}_{w})[i])$checkmark
@@
   HONEST Counterfactual tuple $f(\hat{\mathbb{Y}})=\tfrac{\Sigma_{\hat{Y}_{k}\in\hat{\mathbb{Y}}_{k}}\Sigma_{\hat{y}\in\hat{Y}_{k}}\mathbb{I}_{\textrm{HurtLex}}(\hat{y})}{|\hat{\mathbb{Y}}|\cdot k}$$\times$
   Psycholinguistic Norms Any prompt $f(\hat{\mathbb{Y}})=\tfrac{\Sigma_{\hat{Y}\in\hat{\mathbb{Y}}}\Sigma_{\hat{y}\in\hat{Y}}\textrm{sign}(\textrm{affect-score}(\hat{y}))\textrm{affect-score}(\hat{y})^{2}}{\Sigma_{\hat{Y}\in\hat{\mathbb{Y}}}\sum_{\hat{y}\in\hat{Y}}|\textrm{affect-score}(\hat{y})|}$checkmark
   Gender Polarity Any prompt $f(\hat{\mathbb{Y}})=\tfrac{\Sigma_{\hat{Y}\in\hat{\mathbb{Y}}}\Sigma_{\hat{y}\in\hat{Y}}\textrm{sign}(\textrm{bias-score}(\hat{y}))\textrm{bias-score}(\hat{y})^{2}}{\Sigma_{\hat{Y}\in\hat{\mathbb{Y}}}\sum_{\hat{y}\in\hat{Y}}|\textrm{bias-score}(\hat{y})|}$checkmark
*Data structure corresponds with the task. For example, prompts indicate text generation. $\textrm{}^{\dagger}$tatic word embeddings are not used with LLMs, but we include the word embedding metric WEAT for completeness given its relevance to sentence embedding metrics. $\textrm{}^{\ddagger}$ee § 3.3.1 for definition of $s(\cdot)$.